James earns $5.20 per hour for his basic 40-hour week. When he works overtime he is paid his basic rate plus $0.80. One week he works 'y' hours and receives $298.00. How many hours overtime did he work?
Answer by rfer(16322)

5.20*40=208
298-208=$90 OT
90/(5.20+0.80)=15 hrs OT